OpenStudy (anonymous):

What is the value of k? 6 × 4 × 13 = k × 13 A. 10 B. 24 C. 52 D. 78

OpenStudy (jtvatsim):

Do you have any ideas? What would you think to do first? :)

OpenStudy (anonymous):

multiply 6x4 then multiply the proudut to13

OpenStudy (jtvatsim):

Sounds good! Let's try that and see what happens.

OpenStudy (jtvatsim):

Looks like I'm getting 312 after doing that. How about you?

OpenStudy (anonymous):

yeps

OpenStudy (jtvatsim):

Great, so I guess we are at 312 = k x 13

OpenStudy (jtvatsim):

Now what would you do?

OpenStudy (kendricklamar2014):

Step 1: Simplify both sides:\[6 × 4 × 13 = k × 13 = 312=13 \times k\]

OpenStudy (anonymous):

all i need to know is the value k

OpenStudy (kendricklamar2014):

Step 2: Flip the Equation:\[312=13\times k~becomes~13\times k=312\]

OpenStudy (jtvatsim):

Right! So, we know what 13 times k is, how could you find k by itself? :)

OpenStudy (anonymous):

:D

OpenStudy (kendricklamar2014):

Step 3: Divide both sides by 13:\[\frac{ 13\times k }{ 13 } = \frac{ 312 }{ 13 }\]

OpenStudy (kendricklamar2014):

\[k = 24\]

OpenStudy (jtvatsim):

Hey, look! There's a great idea right there. :) Just divide both sides by 13. This leaves us with k.

OpenStudy (anonymous):

COOLZ

OpenStudy (jtvatsim):

lol Not too bad. Good job!

OpenStudy (anonymous):

THANKS all i need is a push

OpenStudy (jtvatsim):

Yep, from what you started saying in the beginning, it seems like you have the basic ideas in place. With some more practice, you'll be all set.

OpenStudy (anonymous):

:D :D :D